# Polyvalent Machine-Learned Potential for Cobalt: from Bulk to Nanoparticles

### v1

@@ III Application to nanoparticles
In the following analysis, our objective was to assert the accuracy and robustness of the present q-SNAP in describing the structure and energetics of nanoparticles. This capability holds substantial importance, particularly in the context of catalysis, where accurate predictions are crucial for understanding and optimizing catalytic processes.
@@ III.1 Morphological stability at T=0 K
Figure 6: (Color online) Stability of nanoparticles of different shapes as a function of their size at 0 K.
@@
$\Delta(N)=\frac{E(N)-N\epsilon_{coh}}{N^{2/3}},$ (9)
where, $E(N)$ is the total energy of a nanoparticle containing $N$ atoms, and $\epsilon_{coh}$ the cohesive energy of bulk hcp cobalt, which is taken to be -7.04 eV.atom-1.
Farkaš and de Leeuw [26] used DFT-PBE to minimize nanoparticles of different shapes up to 1000 atoms. They determined the relative stabilities of larger systems using a linear extrapolation. Their predictions are comparable to the results of the present explicit calculations, but there are noticeable differences: Farkaš and de Leeuw [26] predict the icosahedron to be the most stable shape up to 5500 atoms, followed by the hcp structure for larger nanoparticles. The present work predicts the transition at 3880 atoms. In Ref. [26] the transition between the truncated octahedron and hcp occurs near 500 atoms while the present results give 1879 atoms. In Ref. [26] the curves of the decahedron and cuboctahedron are separated by minimal differences in energy, but do not intersect while the explicit q-SNAP computations predict a transition at 1315 atoms, as can be seen from Fig. 6.
@@ III.2 Vacancy formation energy on vertices
Figure 8: (Color online) Hexagonal close-packed nanoparticle (587 atoms) minimized by the q-SNAP (left) and the EAM potential (right). Both q-SNAP and DFT-PBE predict minor deviations relative to the initial state (with one vertex removed from the minimized nanoparticle), wherein atoms surrounding the vacancy tend to move away from it. Conversely, the EAM potential predicts an edge shift to fill the vacancy.
@@
According to both potentials, the vacancy formation energy decreases with an increasing number of atoms. The EAM potential predicts that forming a vacancy on a nanoparticle’s vertex is always favorable, except for very small truncated octahedra and hcp nanoparticles, whereas the q-SNAP predicts this behavior only for hcp nanoparticles containing more than 323 atoms, as shown in Fig. 7. For all sizes and shapes, the present q-SNAP predicts that the atoms surrounding the vacancy leave more space for it, while the EAM potential predicts the surrounding atoms to fill the vacancy. For hcp nanoparticles up to 967 atoms, the EAM potential even predicts that the entire edge is shifted to fill the vacancy, as illustrated on the right hand side of Fig. 8. For larger sizes, the edge does not shift, as predicted on the whole range according to the q-SNAP and for the hcp nanoparticle of 153 atoms according to DFT-PBE. The two types of minimizations are represented by two distinct curves in Fig. 7. In addition, Fig. 9 illustrates the excellent agreement between the q-SNAP and DFT-PBE predictions for small nanoparticles.
III.3 Melting
@@
$T_{m,NP}=T_{m,b}\left(1-\frac{C}{D}\right),$ (11)
@@
$T_{m,\text{NP}}=T_{m,\text{bulk}}-aN^{-\frac{1}{3}},$ (12)
@@
Above 693 K, bulk cobalt adopts a face-centered cubic structure. Therefore, to avoid shape effects and to be consistent with the determination of the melting temperature using the biphased cell method, the nanoparticles used in this study are perfect truncated octahedra containing 1289, 2406, 4033, 6266, and 9201 atoms, which correspond to the magic numbers of that shape. These nanoparticles are equilibrated at 900 K, then heated up to 2000 K with an overall heating rate of 0.08 K.ps-1, in intervals of 5 K at constant temperature lasting 62.5 ps each. For these calculations, the system evolves through Langevin dynamics. The melting temperature corresponds to the maximum of the first derivative of the energy with respect to temperature, which was well defined in all cases, as shown in Fig. 10.
@@ IV Conclusion
In this article we presented the development, validation, and capabilities of a q-SNAP to predict static and dynamic properties of bulk Co, its surfaces, and nanoparticles.

### v2

@@ III Application to nanoparticles
In the following analysis, our objective was to exploit the accuracy and robustness of the present q-SNAP in describing the structure and energetics of nanoparticles. This capability holds substantial importance, particularly in the context of catalysis, where accurate predictions are crucial for understanding and optimizing catalytic processes.
@@ III.1 Morphological stability at T=0 K
Figure 7: (Color online) Relative stability of nanoparticles of different shapes as a function of their size at 0 K. The most stable nanoparticles have the lowest $\Delta$ value.
@@
$\Delta(N)=\frac{E(N)-N\epsilon_{coh}}{N^{2/3}}$ (3)
where $E(N)$ is the total energy of a nanoparticle containing $N$ atoms, and $\epsilon_{coh}$ the cohesive energy of bulk hcp cobalt, which is -7.04 eV.atom-1.
Farkaš and de Leeuw [31] used DFT-PBE to minimize nanoparticles of different shapes up to 1000 atoms. They determined the relative stabilities of larger systems using a linear extrapolation. Their predictions are comparable to the results of the present explicit calculations, but there are noticeable differences: Farkaš and de Leeuw [31] predict the icosahedron to be the most stable shape up to 5500 atoms, followed by the hcp structure for larger nanoparticles. The present work, without any extrapolation, predicts the transition at 3880 atoms. In Ref. [31], the transition between the truncated octahedron and hcp occurs near 500 atoms while the present results give 1879 atoms. In Ref. [31], the curves of the decahedron and cuboctahedron are separated by minimal differences in energy, but do not intersect, while the explicit q-SNAP computations predict a transition at 1315 atoms, as can be seen from Fig. 7.
@@ III.2 Vacancy formation energy on vertices
Figure 9: (Color online) Hexagonal close-packed nanoparticle (587 atoms) minimized using DFT-PBE (left), the q-SNAP (middle) and the EAM potential (right). Both q-SNAP and DFT-PBE predict minor deviations relative to the initial state (with one vertex removed from the minimized nanoparticle), wherein atoms surrounding the vacancy tend to move away from it. Conversely, the EAM potential predicts an edge shift to fill the vacancy.
@@
According to both potentials, the vacancy formation energy decreases with an increasing number of atoms, as shown in Fig. 8. The EAM potential predicts that forming a vacancy on a nanoparticle’s vertex is always favorable, except for very small truncated octahedra and hcp nanoparticles, whereas the q-SNAP predicts this behavior only for hcp nanoparticles containing more than 323 atoms. For all sizes and shapes, the present q-SNAP predicts that the atoms surrounding the vacancy leave more space for it, while the EAM potential predicts the surrounding atoms to fill the vacancy. For hcp nanoparticles up to 967 atoms, the EAM potential even predicts that the entire edge is shifted to fill the vacancy, as illustrated on the right hand side of Fig. 9. For larger sizes, the edge does not shift, as predicted on the whole range according to the q-SNAP and for the hcp nanoparticle of 153 atoms according to DFT-PBE. The two types of minimizations are represented by two distinct curves in Fig. 8. In addition, Fig. 10 illustrates the excellent agreement between the q-SNAP and DFT-PBE predictions for small nanoparticles.
III.3 Dynamics
III.3.1 Stability over long simulation times
Figure 11: (Color online) Heat capacity of a 2406 atoms truncated octahedron during a 100 ns simulation at 300 K. The dashed line corresponds to the evolution of its RMSE over time.
Figure 12: (Color online) Truncated octahedron of 2406 atoms at the beginning (left), after 50 ns (middle) and at the end (right) of a 100 ns simulation at 300 K.
The main purpose of this potential is to model large complex systems over timescales that are inaccessible with DFT.
MLPs, particularly Neural Network Potentials (NNPs), are known to become unstable or inaccurate after a certain simulation time [69]. To test the stability of the new q-SNAP, we ran a simulation of 100 ns on a 2406 atoms truncated octahedron. The nanoparticle was placed in a cubic cell with 50 Å of vacuum around it and evolved through Langevin dynamics at a constant temperature of 300 K, with a timestep of 2 fs. The instantaneous total energy of the system is written every 20 fs.
From these data, the heat capacity is computed as
$C_{V}=\frac{\langle E^{2}\rangle-\langle E\rangle^{2}}{N_{A}nk_{B}T^{2}}$ (5)
where $E$ is the total energy of the system, $N_{A}$ Avogadro’s number, $n$ the number of atoms in the nanoparticle (here 2406), $k_{B}$ the Boltzmann’s constant and $T$ the temperature (here 300 K). The heat capacity, resulting from the average of the values computed between 50 and 100 ns is 25.14 J.K-1.mol-1. This value is 9.4% larger than that for bulk hcp Co (see Table 3). The larger specific heat of nanoparticles is due to the lower frequency of vibrational modes of surface atoms compared to bulk atoms [70].
Fig. 11 shows its constancy throughout the entire simulation, as well as its RMSE, which decreases over time, as expected. Additionally, Fig. 12 shows the nanoparticle at the beginning, in the middle, and at the end of the simulation, always maintaining the same shape. The remarkable stability of the presented q-SNAP is a key strength.
III.3.2 Melting
@@
$T_{m,NP}=T_{m,b}\left(1-\frac{C}{D}\right)$ (6)
@@
$T_{m,\text{NP}}=T_{m,\text{bulk}}-aN^{-\frac{1}{3}}$ (7)
@@
Above 693 K, bulk cobalt adopts a face-centered cubic structure. Therefore, to avoid shape effects and to be consistent with the determination of the melting temperature using the biphased cell method, the nanoparticles used in this study are perfect truncated octahedra containing 1289, 2406, 4033, 6266, and 9201 atoms, which correspond to the magic numbers of that shape. These nanoparticles are equilibrated at 900 K, then heated up to 2000 K for 13.75 ns. We chose to split the trajectory in intervals of 62.5 ps at constant temperature, every 5 K, to be sure that the energy is converged. Thus, the overall heating rate is 0.08 K.ps-1. For these calculations, the system evolves through Langevin dynamics. The melting temperature corresponds to the maximum of the first derivative of the energy with respect to temperature, which was well defined in all cases, as shown in Fig. 13.
@@ IV Conclusion
In this article we presented the development, validation, and application of a q-SNAP to predict static and dynamic properties of bulk Co, surfaces, and nanoparticles.